OpenStudy (bemoseemo):

The length of a rectangle is one unit shorter than one-sixth of the width, x. Enter a simplified expression that represents the perimeter of the rectangle. Note: Use one variable and a fraction in the answer.

OpenStudy (bemoseemo):

its kinda confusing when you look at it

OpenStudy (3mar):

Not at all. It will be simple now. I will show you.

OpenStudy (bemoseemo):

I know that l-1=w*1/6

OpenStudy (3mar):



OpenStudy (3mar):

\[langth=1-\frac{ 1 }{ 6 }x\] now you have the length and the width, can you express the perimeter of the rectangular?

OpenStudy (3mar):

one-sixth=1/6 shorter than=1/6- unit shorter than one-sixth=1/6 - 1

OpenStudy (3mar):

got it?

OpenStudy (bemoseemo):

yeah i think so ty so much

OpenStudy (3mar):

I told you that "It will be simple now".

OpenStudy (3mar):

need more help?

OpenStudy (3mar):

Thank you for the medal!